OpenStudy (anonymous):

Help on Math Investigation about square numbers? Do you have any ideas or suggestions?

OpenStudy (anonymous):

square number???

OpenStudy (anonymous):

like 9, 16, etc...

OpenStudy (anonymous):

ok ok when a number is multiplied by its self one time only the resulting number is called the square of that number.

OpenStudy (anonymous):

all square numbers numbers can be written as the factors of other numbers

OpenStudy (anonymous):

but those factors are always the same

OpenStudy (anonymous):

i meant discover new concepts about square numbers. we're supposed to do an investigation about it but we have no ideas so far.

OpenStudy (anonymous):

ok i have one idea

OpenStudy (anonymous):

this idea is intresting but you have to give me some time its really intresting

OpenStudy (anonymous):

ok. sure what is your idea?

OpenStudy (anonymous):

consider a square number let say 4 square it again you got 16 now take the consecutive numbers before and after 4 that are 3 and 5 multiply them we get 15 now 16 - 15 = 1 this result is the same for all square numbers

OpenStudy (anonymous):

@danielle02

OpenStudy (anonymous):

@03453660 that's awesome! thanks for the help :)

OpenStudy (anonymous):

take any square number do the process i did you will always get 1

OpenStudy (anonymous):

another idea: a sequence of odd numbers is generated by taking the difference of consecutive square numbers

OpenStudy (anonymous):

like 1-0 = 1 4-1=3 9-4=5 16 - 9 = 7 ............ .......... ........

OpenStudy (anonymous):

are these ideas new or not?????

OpenStudy (anonymous):

that's a good one too. i like the first one more.

OpenStudy (anonymous):

yeah, thank you very much for your help! i'll try to come up with new ideas based on your help :) @03453660

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

by the way i like your question too much

OpenStudy (anonymous):

@danielle02 did you find anything more about square numbers

OpenStudy (tyteen4a03):

Consider the perfect square identity: (x+y)^2 = x^2 + 2xy + y^2, and difference of squares identity; x^2 - y^2 = (x+y)(x-y)

OpenStudy (anonymous):

@03453660 i'm still investigating. i'm currently trying to find a pattern so i could come up with something.

OpenStudy (anonymous):

@tyteen4a03 what do you mean by this?

OpenStudy (tyteen4a03):

@danielle02 These identity are related to square numbers. You can google the proof of these identities.

OpenStudy (anonymous):

isn,t the pattern of odd numbers i told you before was enough???

OpenStudy (anonymous):

i'm trying to find more patterns you know, i do have fun investigating square numbers because they're interesting

OpenStudy (anonymous):

exactly i didnt thought over these numbers but now i found these numbers very intresting